OpenStudy (anonymous):

write an equation in slope intercept form for the line described. passes (-2,3) and (0,1)

OpenStudy (anonymous):

y=-x+1

OpenStudy (anonymous):

I came up with y= -3/2x +1 m=-3/2, b=1 not sure what I did wrong

OpenStudy (anonymous):

m=-1

OpenStudy (anonymous):

\[m=\frac{3-1}{-2-0}=-1\]

OpenStudy (anonymous):

OK I see and then the b=1 is that correct ?